OpenStudy (anonymous):

any help again? just asking for double check. :) Two kinds of coffee are blended to produce a mixture of 25 kilograms worth $52 a kilogram. The costs of the two kinds of coffee are $40 and $56 a kilogram. Find how many kilograms of each kind are blended. I also need the solution. THANKS! :)

OpenStudy (anonymous):

the equation you want to solve if [40x+56(25-x)=25\times 52\]

OpenStudy (anonymous):

lets try that again. \[40x+56(25-x)=25\times 52\]

OpenStudy (anonymous):

so x will be the kilogram of each kind?

OpenStudy (anonymous):

reason is if you put x = amount of $40 beans then it cost 40x . if you use x of one kind and the total is 25 then you use 25-x of the other kind that cost 56, for total of 56(25-x) and finally you end up with 25 kilos at 52 for a total of 25*52

OpenStudy (anonymous):

no the way i wrote it x will be the kilos of $40 beans

OpenStudy (anonymous):

and 25-x will be the kilos of $56 beans

OpenStudy (anonymous):

okay. got that. then?

OpenStudy (anonymous):

we solve \[40x+56(25-x)=52\times 25\] \[40x+1400-56x-1300\] \[-16x=1300-1400\] \[-16x=-100\] \[x=\frac{100}{16}\] \[x=6\tfrac{1}{4}\]

OpenStudy (anonymous):

you should check my algebra because it is late. i may have made a mistake but i do not see it

OpenStudy (anonymous):

okay. :)

OpenStudy (anonymous):

so i get \[6\tfrac{1}{4}\] of the $40 beans and \[25-6\tfrac{1}{4}=18\tfrac{3}{4}\] of the $56 beans

OpenStudy (anonymous):

thanks for the help! :)